OpenStudy (swissgirl):

What is the fourth derivative of \(f(x)=3xe^x-e^{2x}\) ?

OpenStudy (bahrom7893):

well f'(x*e^x) = x*e^x + e^x = e^x(x+1). So: f'(x) = 3e^x(x+1) - 2e^(2x) f''(x) = (3xe^x+3e^x - 2e^(2x))' = 3e^x(x+1)+3e^x - 4e^(2x)

OpenStudy (bahrom7893):

f'''(x) = (3xe^x + 3e^x + 3e^x - 4e^(2x))' = (3xe^x + 6e^x - 4e^(2x))' = 3e^x(x+1) + 6e^x - 8e^(2x)

OpenStudy (bahrom7893):

I think I made a mistake in f''(x)

hero (hero):

Yes, you did

OpenStudy (bahrom7893):

No i didn't

hero (hero):

You made a mistake somewhere

OpenStudy (bahrom7893):

f''''(x) = (3xe^x+3e^x+6e^x-8e^(2x))' = (3xe^x + 9e^x - 8e^(2x))' = 3e^x(x+1)+9e^x - 16e^(2x) <-Final answer

OpenStudy (swissgirl):

Btw u didnt go wrong

OpenStudy (swissgirl):

Thankkksss Bahhhrrrooommmmmmm

OpenStudy (bahrom7893):

Hero.. like i said.. ignore me.

hero (hero):

Yeah, that's the correct answer

hero (hero):

Why do you want me to ignore you @bahrom7893